# Stop treating ‘AGI’ as the north-star goal of AI research

### v3

@@ Abstract
The AI research community plays a vital role in shaping the scientific, engineering, and societal goals of AI research. In this position paper, we argue that focusing on the highly contested topic of ‘artificial general intelligence’ (‘AGI’) undermines our ability to choose effective goals. We identify six key traps—obstacles to productive goal setting—that are aggravated by AGI discourse: Illusion of Consensus, Supercharging Bad Science, Presuming Value-Neutrality, Goal Lottery, Generality Debt, and Normalized Exclusion. To avoid these traps, we argue that the AI research community needs to (1) prioritize specificity in engineering and societal goals, (2) center pluralism about multiple worthwhile approaches to multiple valuable goals, and (3) foster innovation through greater inclusion of disciplines and communities. Therefore, the AI research community needs to stop treating “AGI” as the north-star goal of AI research.
@@ 1 Introduction
Recent advances in large language models (LLMs) have sparked interest in “achieving human-level ‘intelligence”’ as a “north-star goal” of the AI field (Morris et al., 2023; McCarthy et al., 1955). This goal is often referred to as “artificial general intelligence” (“AGI”) (Chollet, 2024a; Tibebu, 2025). Yet rather than helping the field converge around shared goals, AGI discourse has mired it in controversies. Researchers diverge on what AGI is and assumptions about goals and risks (Summerfield, 2023; Morris et al., 2023; Blili-Hamelin et al., 2024). Researchers further contest the motivations, incentives, values, and scientific standing of claims about AGI (Gebru & Torres, 2024; Mitchell, 2024; Ahmed et al., 2024; Altmeyer et al., 2024). Finally, the building blocks of AGI as a concept—intelligence and generality—are contested in their own right (Gould, 1981; Anderson, 2002; Hernandez-Orallo & O’hEigeartiagh, 2018; Cave, 2020; Raji et al., 2021; Alexandrova & Fabian, 2022; Blili-Hamelin & Hancox-Li, 2023; Hao, 2023; Guest & Martin, 2024; Paolo et al., 2024; Mueller, 2024).
@@
Building on prior work on the ambiguity between exploratory and confirmatory research in ML (Herrmann et al., 2024), unscientific performance claims (Altmeyer et al., 2024), SOTA-chasing (Raji et al., 2021), homogenization of research approaches (Kleinberg & Raghavan, 2021; Fishman & Hancox-Li, 2022; Bommasani et al., 2022), the values embedded in ML research (Birhane et al., 2022b), and more, our account identifies key obstacles to productive goal setting in AI research—traps. [^1] Our terminology parallels Selbst et al. (2019) on fairness. We argue that the AGI narrative makes it more difficult to overcome each obstacle. To avoid these traps, we posit that communities should stop treating AGI as a north-star goal of AI research.
@@
An overarching theme in our discussion is the research community’s unique responsibility to help distinguish hype from reality. The outputs of AI research are deployed as real-world products at a staggering pace, in proliferating contexts, affecting billions of people. This warrants urgent work on trusted, evidence-based answers to questions about the scientific, engineering, and societal merits of AI tools. As argued by the U.N.’s AI Advisory Body, there is “an overwhelming amount of information… making it difficult to decipher hype from reality. This can fuel confusion, forestall common understanding and advantage major AI companies at the expense of policymakers, civil society and the public,” (United Nations, 2024). Our position paper addresses this theme: Each of the 6 traps in our account is an obstacle to distinguishing hype from reality.
@@
A secondary theme in our discussion is the relationship between people and technology. Ultimately, we argue that instead of a single north-star goal, the AI community needs to pursue multiple specific scientific, engineering, and societal goals. Narratives about the future of complex technologies are, we argue, prone to blurring hype and reality. If building consensus around an alternative unifying goal proves useful, we propose the goal of supporting and benefiting human beings.
@@
In the next section, we examine six ‘traps’ in AI research—obstacles to productive goal setting (§2). We argue that AGI discourse reinforces and amplifies each problem. Subsequently, we defend three recommendations to avoid the traps (§3): specificity of goals; pluralism of goals and approaches; and more inclusive goal setting. Our concluding rebuttal champions our position against an alternative view4): that AGI should remain the north-star goal of the field.
Because the contested nature of AGI is a central theme in the present paper, we avoid providing our own definition for the term. Instead, we provide example definitions throughout the present discussion, as well as a table of illustrative definitions (Appendix A).
@@ 2 Traps
Although our discussion strongly focuses on AGI, we believe that our points apply to the broader project of developing trusted, evidence-based approaches to deliberating over the scientific, engineering, and societal goals of AI research.
The problems we discuss are highly interrelated. For instance, SOTA-chasing, discussed in relationship to the role of misaligned incentives in shaping goal setting (§2.4), also has implications for bad science (§2.2). Similarly, the problem of the lack of consensus on AGI (§2.1), is a theme that recurs throughout. We do not intend the traps to be mutually exclusive. Rather, our goal for each trap is to provide distinct and useful resources for mitigating failure modes in productive goal setting.
@@ 2.1 Illusion of Consensus
Using shared term(s) in a way that gives a false impression of consensus about goals, despite goals being contested.
The increasing popular use of the term “AGI” (Holland, 2025; Grossman, 2023; IBM, 2023) creates a sense of familiarity, giving the illusion that there is a shared understanding on what AGI is, and broad agreement on research goals in AGI development. However, there are vastly different opinions on what the term AGI refers to, what an AGI research agenda looks like, and what the goals in AGI development are. Left unchecked, this illusion obstructs explicit engagement on what the goals of AI research are and should be.
From popular discourse to research papers to corporate marketing materials, the vast majority of references to AGI fall into this trap when they uncritically cite claims about so-called AGI. For examples of uncritical media claims, see Grossman (2023); IBM (2023); see Altmeyer et al. (2024) for examples of overhyped research. [^2] Some researchers who advocate for AGI as a goal have avoided the Illusion of Consensus trap (§2.1); e.g., Morris et al. (2023) explicitly call for investigating disagreements about goals, predictions, and risks that underpin prominent accounts of AGI. Summerfield (2023) summarizes the issue: “AI researchers hope to discover how to build AGI. The problem is that nobody really knows exactly what an AGI would look like.” Mueller (2024), in turn, calls AGI “a meaningless concept, an emperor with no clothes.” Blili-Hamelin et al. (2024) identify multiple types of disagreement among attempts to define AGI or human-level AI. The contested nature of AGI as a goal becomes even more acute in critiques of AGI concepts (see, e.g., Altmeyer et al., 2024; Mueller, 2024; Van Rooij et al., 2024).
Beyond AGI, AI research is rife with topics that involve disagreement about goals, values, and concepts. For example, Mulligan et al. (2016) argue that privacy should be understood as an “essentially contested concept.” They argue lack of agreement about the meaning and significance of privacy is not merely a matter of confusion—rather, disagreement and contestation are desirable features that enable privacy to adapt to changing technical and social contexts. Similarly, there is now widespread understanding fairness should be understood as a contested topic, not only admitting incompatible mathematical formalizations but also incompatible values, worldviews, and theoretical assumptions (Friedler et al., 2021; Jacobs & Wallach, 2021).
@@ 2.2 Supercharging Bad Science
Poorly defined concepts and experimental procedures deployed in pursuit of AGI worsen current problems with bad science in AI.
@@
Research that produces reliable empirical knowledge about AI is vital to public interest decisions about AI’s potential for societal and environmental benefit and harm. Yet, many experts have noted a pervasive lack of scientific grounding in AI research (United Nations, 2024; Sloane et al., 2022; Narayanan & Kapoor, 2024; Raji et al., 2022; Widder & Hicks, 2024; Suchman, 2023; Hullman et al., 2022; Van Rooij et al., 2024; Guest & Martin, 2024). We argue that vagueness in AGI discourse exacerbates existing problems with the scientific validity of AI research.
@@
Problem 1: Underspecification and external validity. One problem with the pursuit of AGI as a concrete goal is underspecification, where lack of specificity in goals or concepts leads to cascading epistemic problems, including irrefutability, lack of external validity, flawed experimental design, and flawed evaluation. These common problems in AI research are worsened in the AGI context by the lack of scientifically grounded definitions of AGI (§2.1).
@@
Underspecification of learning goals also undermines external validity—the question of whether a measurement corresponds to the real-world phenomenon it’s supposed to capture. A good example is debate about whether “language understanding” benchmarks actually measure language understanding (Jacobs & Wallach, 2021; Liao et al., 2021).
@@
External validity is also relevant when researchers equate human faculties with model proxies (Hullman et al., 2022), such as claiming that a model “capable of linking specific objects with more general visual context” is evidence of “imagination” (Fei et al., 2022). This rhetorical move is enabled by using colloquial terms like “imagination” without considering whether it corresponds to the human faculty. Altmeyer et al. (2024) likewise critiques Gurnee & Tegmark (2024) for inflated claims enabled by the vagueness of the term “world model.” Underspecified goals trickle down into many areas of experimental design, such as learning pipelines, evaluation metrics, tasks, representations, and methods.
@@
External validity is also undermined when researchers claim to measure concepts from other fields, like intelligence. The fields of psychology, neuroscience, and cognitive science have studied human intelligence for generations, yet even they lack consensus on what “intelligence” is (Hao, 2023; Gopnik, 2019). Conversely, AI research is no longer concerned with modeling human cognition (Van Rooij et al., 2024; Guest & Martin, 2024). Instead, AI developers define “intelligence” on their own terms, privileging definitions convenient for benchmarking or selling products (§2.4), while benefiting from historically positive connotations of the term “intelligence”.
@@
Problem 2: Ambiguity between science and engineering. Another problem with the pursuit of AGI is confusion between science and engineering (Agre, 2014; Hutchinson et al., 2022; Altmeyer et al., 2024). As Hullman et al. (2022) point out, a “typical supervised ML paper” (one that shows some accuracy metrics on a benchmark) is often just an “engineering artifact”, a tool attached to performance claims that cannot be refuted because of replication challenges. Altmeyer et al. (2024) argue that this ambiguity between science and engineering means rigorous hypothesis testing with “specific conditions and considering effect sizes” is often omitted, with results often presented as “engineering achievements” without specifying precisely what is being tested, relevant hypotheses, and what effect sizes would constitute a null hypothesis.
@@
This ambiguity invites experimenter and confirmation biases, since researchers are incentivized to “pay little or no attention to competing hypotheses or explanations” or “[fail] to articulate a sufficiently strong null hypothesis,” (Altmeyer et al., 2024). Confusion between science and engineering also manifests when it is unclear if a study is pursuing scientific goals—of explanation, hypothesis confirmation, etc.—or goals of specific engineering applications—e.g., a proof-of-concept (Hutchinson et al., 2022). This exacerbates questions about external validity: without clear and specific experimental goals, it is easier to provide post-hoc interpretations of experiments that “support” a wide variety of goals.
@@ 2.3 Presuming Value-Neutrality
Framing goals as purely technical or scientific in nature when political, social, or ethical considerations are implied and/or essential.
Presuming Value-Neutrality occurs when technical or scientific goals become disconnected from their value-laden assumptions: aspects of AI research that are—and should be—informed by political, social, and ethical considerations. The AI research community has recently begun examining these value-laden assumptions (Zhao et al., 2024; Fishman & Hancox-Li, 2022; Dotan & Milli, 2020; Birhane et al., 2022b; Scheuerman et al., 2021; Hutchinson et al., 2022; Bommasani, 2022; Denton et al., 2020, 2021; Mathur et al., 2022; Shilton, 2018; Broussard et al., 2019; Green, 2021; Blodgett et al., 2020; Viljoen, 2021; Abebe et al., 2020; Birhane & Guest, 2021; Blili-Hamelin & Hancox-Li, 2023; Blili-Hamelin et al., 2024; Costanza-Chock, 2020).
When efforts to define AGI and related concepts do not explicitly examine the societal goals and values embedded in their definitions, they fall into the Presuming Value-Neutrality trap. Examples include proposals for “universal intelligence” (Legg & Hutter, 2007b; Hernández-Orallo et al., 2014).
@@ 2.4 Goal Lottery
This trap occurs when incentives, circumstances, or luck drive the adoption of goals inadequately supported by scientific, engineering, or societal merit.
Researchers have studied the role of socioeconomic factors, trends, and circumstantial factors in shaping AI research. For instance, Hooker (2021) has argued that a form of hardware lottery—the greater availability of hardware with strengths in parallel processing—was key to the resurgence of deep learning in the 2010s. [^3] On similar lottery or path dependence effects, see Dehghani et al. (2021); Fishman & Hancox-Li (2022); Hooker (2024); Liebowitz & Margolis (1995); Peacock (2009); Bauer & Gill (2024); Rossbach (2023). Similarly, researchers have examined the role of incentives, socioeconomic factors, and hype cycles in AI research (Narayanan & Kapoor, 2024; Wang et al., 2024; Hicks et al., 2024; Gebru & Torres, 2024; Sartori & Bocca, 2023; Delgado et al., 2023; Raji et al., 2022; Widder & Nafus, 2023). With this trap, we focus on cases where lotteries (luck) or incentives drive the adoption of unjustified goals—goals that are inadequately supported by scientific, engineering, or societal merit.
@@
Another example is benchmark SOTA-chasing—pursuing top scores on popular benchmarks (Bender et al., 2021; Church & Kordoni, 2022; Hullman et al., 2022; Raji et al., 2021). Despite strong professional incentives encouraging this practice, it lacks engineering, scientific, and societal justification. Benchmarks poorly reflect model performance in real application contexts because of problems like data leakage, overfitting to benchmarks, and data heterogeneity (Balloccu et al., 2024; Xu et al., 2024; Zhang et al., 2024; El-Mhamdi et al., 2022; Hanneke & Kpotufe, 2022; El-Mhamdi et al., 2021). In short, the measurement method lacks external validity. Yet the practice persists due to reputational and financial rewards, demonstrating misalignment between incentivized goals and their actual merits.

### v4

# Position: Stop Treating ‘AGI’ as the North-star Goal of AI Research

@@ Abstract
The AI research community plays a vital role in shaping the scientific, engineering, and societal goals of AI research. In this position paper, we argue that focusing on the highly contested topic of ‘artificial general intelligence’ (‘AGI’) undermines our ability to choose effective goals. We identify six key traps—obstacles to productive goal setting—that are aggravated by AGI discourse: Illusion of Consensus, Supercharging Bad Science, Presuming Value-Neutrality, Goal Lottery, Generality Debt, and Normalized Exclusion. To avoid these traps, we argue that the AI research community needs to (1) prioritize specificity in scientific, engineering, and societal goals, (2) center pluralism about multiple worthwhile approaches to multiple valuable goals, and (3) foster innovation through greater inclusion of disciplines and communities. Therefore, the AI research community needs to stop treating “AGI” as the north-star goal of AI research.
@@ 1 Introduction
Recent advances in large language models (LLMs) have sparked interest in “achieving human-level ‘intelligence”’ as a “north-star goal” of the AI field (McCarthy et al., 1955; Morris et al., 2024). This goal is often referred to as “artificial general intelligence” (“AGI”) (Chollet, 2024a; Tibebu, 2025). Yet rather than helping the field converge around shared goals, AGI discourse has mired it in controversies. Researchers diverge on what AGI is and assumptions about goals and risks (Summerfield, 2023; Morris et al., 2024; Blili-Hamelin et al., 2024). Researchers further contest the motivations, incentives, values, and scientific standing of claims about AGI (Gebru & Torres, 2024; Mitchell, 2024; Ahmed et al., 2024; Altmeyer et al., 2024). Finally, the building blocks of AGI as a concept—intelligence and generality—are contested in their own right (Gould, 1981; Anderson, 2002; Hernández-Orallo & Seán Ó hÉigeartaigh, 2018; Cave, 2020; Raji et al., 2021; Alexandrova & Fabian, 2022; Blili-Hamelin & Hancox-Li, 2023; Hao, 2023; Guest & Martin, 2024; Paolo et al., 2024; Mueller, 2024).
@@
Building on prior work on the ambiguity between exploratory and confirmatory research in ML (Herrmann et al., 2024), unscientific performance claims (Altmeyer et al., 2024), SOTA-chasing (Raji et al., 2021; Church & Kordoni, 2022), homogenization of research approaches (Kleinberg & Raghavan, 2021; Fishman & Hancox-Li, 2022; Bommasani et al., 2022), the values embedded in ML research (Birhane et al., 2022b), and more, our account identifies key obstacles to productive goal setting in AI research—traps. [^1] Our terminology parallels Selbst et al. (2019) on fairness. We provide them here as a diagnosis of problems in goal-setting that we believe are normatively worth addressing, but that the AGI narrative makes difficult to overcome. To avoid these traps, we posit that communities should stop treating AGI as the north-star goal [^2] Sailors who navigate by the astronomical North Star use it to orient their travels toward a desired destination on Earth. With AGI, some researchers are using AGI as a “guiding star” to orient their AI research “travels” towards. Other researchers, however, are actually hoping and working towards the goal of “arriving” at AGI. Our paper argues against both of these approaches. of AI research.
@@
An overarching theme in our discussion is the research community’s unique responsibility to help distinguish hype from reality. The outputs of AI research are deployed as real-world products at a staggering pace, in proliferating contexts, affecting billions of people. This warrants urgent work on trusted, evidence-based answers to questions about the scientific, engineering, and societal merits of AI tools. As argued by the U.N.’s AI Advisory Body, there is “an overwhelming amount of information… making it difficult to decipher hype from reality. This can fuel confusion, forestall common understanding and advantage major AI companies at the expense of policymakers, civil society and the public” (United Nations, 2024). Our position paper addresses this theme: Each of the six traps in our account is an obstacle to distinguishing hype from reality.
@@
A secondary theme in our discussion is the relationship between people and technology. Ultimately, we argue that instead of a single north-star goal, the AI community needs to pursue multiple specific scientific, engineering, and societal goals. If building consensus around an alternative unifying goal proves useful, we propose the goal of supporting and benefiting human beings.
@@
In the next section, we examine six ‘traps’ in AI research—obstacles to productive goal setting (§2). We argue that AGI discourse reinforces and amplifies each problem. Subsequently, we provide three recommendations for avoiding these traps (§3): specificity of goals; pluralism of goals and approaches; and more inclusive goal setting. We conclude by offering a rebuttal against an alternative view4): that AGI should remain the north-star goal of the field.
Because the contested nature of AGI is a central theme in the present paper, we avoid providing our own definition for the term. Instead, we provide example definitions throughout the present discussion, as well as a table of illustrative definitions in Appendix A. [^3] Arguably, many of the concerns we raise about AGI apply to other terms used to refer to future forms of AI, such as “powerful AI” (Amodei, 2024) and “transformative AI” (Gruetzemacher & Whittlestone, 2022). Ultimately, as we revisit in §4, our account can be viewed as critically interrogating north-star goals more generally. We detail how AGI currently serves as a north-star goal in Appendix B.
@@ 2 Traps
The problems we discuss are highly interrelated. For instance, SOTA-chasing, discussed in relationship to the role of misaligned incentives in shaping goal setting (§2.4), also has implications for bad science (§2.2). Similarly, the problem of the lack of consensus about AGI (§2.1) is a theme that recurs throughout. We do not intend the traps to be mutually exclusive. Rather, our goal for each trap is to provide distinct and useful insights for mitigating failure modes in productive goal setting.
@@ 2.1 Illusion of Consensus
Using shared term(s) in a way that gives a false impression of consensus about goals, despite goals being contested
The popular use of the term “AGI” (Grossman, 2023; IBM, 2023; Holland, 2025) creates a sense of familiarity, giving the illusion that there is a shared understanding on what AGI is, and broad agreement on research goals in AGI development. However, there are vastly different opinions on what the term AGI refers to, what an AGI research agenda looks like, and what the goals in AGI development are. Left unchecked, this illusion obstructs explicit engagement on what the goals of AI research are and should be.
From popular discourse to research papers to corporate marketing materials, the vast majority of references to AGI fall into this trap when they uncritically cite claims about so-called AGI. For examples of uncritical media claims, see Grossman (2023) and IBM (2023); see Altmeyer et al. (2024) for examples of overhyped research. [^4] Some researchers who advocate for AGI as a goal have avoided the Illusion of Consensus trap (§2.1); e.g., Morris et al. (2024) explicitly call for investigating disagreements about goals, predictions, and risks that underpin prominent accounts of AGI. Summerfield (2023) summarizes the issue: “AI researchers hope to discover how to build AGI. The problem is that nobody really knows exactly what an AGI would look like.” Mueller (2024) calls AGI “a meaningless concept, an emperor with no clothes.” Blili-Hamelin et al. (2024) identify multiple types of disagreement among definitions of AGI or human-level AI. The contested nature of AGI as a goal is even more acute in critiques of AGI concepts (e.g., Altmeyer et al., 2024; Mueller, 2024; Van Rooij et al., 2024).
Beyond AGI, AI research is rife with topics that involve disagreement about goals, values, and concepts. For example, Mulligan et al. (2016) argue that privacy should be understood as an “essentially contested concept.” They argue lack of agreement about the meaning and significance of privacy is not merely a matter of confusion—rather, disagreement and contestation are desirable features that enable privacy to adapt to changing technical and social contexts. Similarly, there is now widespread acceptance fairness should be understood as a contested topic, not only admitting incompatible mathematical formalizations but also incompatible values, worldviews, and theoretical assumptions (Friedler et al., 2021; Jacobs & Wallach, 2021).
@@ 2.2 Supercharging Bad Science
Worsening current problems with bad science in AI due to poorly defined concepts and experimental procedures
@@
Research that produces reliable empirical knowledge about AI is vital to public interest decisions about AI’s potential for societal and environmental benefit and harm. Yet, many experts have noted a pervasive lack of scientific grounding in AI research (Hullman et al., 2022; Raji et al., 2022; Sloane et al., 2022; Suchman, 2023; Guest & Martin, 2024; Narayanan & Kapoor, 2024; United Nations, 2024; Van Rooij et al., 2024; Widder & Hicks, 2024). We argue that vagueness in AGI discourse exacerbates existing problems with the scientific validity of AI research.
@@
Problem 1: Underspecification and external validity. One problem with the pursuit of AGI as a concrete goal is underspecification (D’Amour et al., 2022), where lack of specificity in goals or concepts leads to cascading epistemic problems, including irrefutability, lack of external validity, flawed experimental design, and flawed evaluation. These common problems in AI research are worsened in the AGI context by the lack of scientifically grounded definitions of AGI (§2.1).
@@
Underspecification of learning goals also undermines external validity—the question of whether a measurement corresponds to the real-world phenomenon it’s supposed to capture. A good example is the debate about whether “language understanding” benchmarks actually measure language understanding (Jacobs & Wallach, 2021; Liao et al., 2021).
@@
External validity is also relevant when researchers equate human faculties with model proxies (Hullman et al., 2022), such as claiming that a model “capable of linking specific objects with more general visual context” is evidence of “imagination” (Fei et al., 2022). This rhetorical move is enabled by using colloquial terms like “imagination” without considering whether it corresponds to the human faculty. Altmeyer et al. (2024) likewise critiques Gurnee & Tegmark (2024) for inflated claims enabled by the vagueness of the term “world model”. Underspecified goals trickle down into many areas of experimental design, such as learning pipelines, evaluation metrics, tasks, representations, and methods.
@@
External validity is also undermined when researchers claim to measure concepts from other fields, like intelligence. The fields of psychology, neuroscience, and cognitive science have studied human intelligence for generations, yet even they lack consensus on what “intelligence” is (Gopnik, 2019; Hao, 2023). Conversely, AI research is no longer concerned with modeling human cognition (Guest & Martin, 2024; Van Rooij et al., 2024). Instead, AI developers define “intelligence” on their own terms, privileging definitions convenient for benchmarking or selling products (§2.4), while benefiting from historically positive connotations of the term “intelligence”.
@@
Problem 2: Ambiguity between science and engineering. Another problem with the pursuit of AGI is confusion between science and engineering (Agre, 2014; Hutchinson et al., 2022; Altmeyer et al., 2024). As Hullman et al. (2022) point out, a “typical supervised ML paper” (e.g., one that reports accuracy metrics on a benchmark) is often just an “engineering artifact”, a tool attached to performance claims that cannot be refuted because of replication challenges. Altmeyer et al. (2024) argue that this ambiguity between science and engineering means rigorous hypothesis testing with “specific conditions and considering effect sizes” is often omitted, with results often presented as “engineering achievements” without specifying precisely what is being tested, relevant hypotheses, and what effect sizes would constitute substantial findings.
@@
This ambiguity invites experimenter and confirmation biases, since researchers are incentivized to “pay little or no attention to competing hypotheses or explanations” or “[fail] to articulate a sufficiently strong null hypothesis,” (Altmeyer et al., 2024). Confusion between science and engineering also manifests when it is unclear if a study is pursuing scientific goals—of explanation, hypothesis confirmation, etc.—or goals of specific engineering applications—e.g., a proof-of-concept (Hutchinson et al., 2022). This exacerbates questions about external validity: without clear and specific experimental goals, it is easier to provide post-hoc interpretations of experiments that “support” a wide variety of goals (§2.4).
@@ 2.3 Presuming Value-Neutrality
Framing goals as purely technical or scientific, when they are in fact laden with political, social, or ethical values
Presuming Value-Neutrality occurs when technical or scientific goals become disconnected from their value-laden assumptions: aspects of AI research that are—and should be—informed by political, social, and ethical considerations. The AI research community has recently begun examining these value-laden assumptions (Shilton, 2018; Broussard et al., 2019; Abebe et al., 2020; Blodgett et al., 2020; Costanza-Chock, 2020; Denton et al., 2020, 2021; Dotan & Milli, 2020; Birhane & Guest, 2021; Green, 2021; Scheuerman et al., 2021; Viljoen, 2021; Birhane et al., 2022b; Bommasani, 2023; Fishman & Hancox-Li, 2022; Hutchinson et al., 2022; Mathur et al., 2022; Blili-Hamelin & Hancox-Li, 2023; Blili-Hamelin et al., 2024; Zhao et al., 2024).
When efforts to define AGI and related concepts do not explicitly examine the societal goals and values embedded in their definitions, they fall into the Presuming Value-Neutrality trap. Examples include proposals for “universal intelligence” (Legg & Hutter, 2007; Hernández-Orallo et al., 2014).
@@ 2.4 Goal Lottery
Adopting goals which are not adequately justified by scientific, engineering, or social merit, but instead on the basis of incentives, circumstances, or luck
Researchers have studied the role of socioeconomic factors, trends, and circumstantial factors in shaping AI research. For instance, Hooker (2021) has argued that a form of hardware lottery—the greater availability of hardware with strengths in parallel processing—was key to the resurgence of deep learning in the 2010s. [^5] On similar lottery or path dependence effects, see Liebowitz & Margolis (1995); Peacock (2009); Dehghani et al. (2021); Fishman & Hancox-Li (2022); Rossbach (2023); Bauer & Gill (2024); Hooker (2024). Similarly, researchers have examined the role of incentives, socioeconomic factors, and hype cycles in AI research (Raji et al., 2022; Delgado et al., 2023; Sartori & Bocca, 2023; Widder & Nafus, 2023; Gebru & Torres, 2024; Hicks et al., 2024; Narayanan & Kapoor, 2024; Wang et al., 2024). With this trap, we focus on cases where lotteries (luck) or incentives drive the adoption of unjustified goals—goals that are inadequately supported by scientific, engineering, or societal merit.
@@
Another example is benchmark SOTA-chasing—pursuing top scores on popular benchmarks (Bender et al., 2021; Raji et al., 2021; Church & Kordoni, 2022; Hullman et al., 2022). Despite strong professional incentives encouraging this practice, it lacks scientific, engineering, and societal justification. Benchmarks poorly reflect model performance in real application contexts because of problems like data leakage, overfitting to benchmarks, and data heterogeneity (El-Mhamdi et al., 2021, 2023; Hanneke & Kpotufe, 2022; Balloccu et al., 2024; Xu et al., 2024; Zhang et al., 2024). In short, the measurement method lacks external validity. Yet the practice persists due to reputational and financial rewards, demonstrating misalignment between incentivized goals and their actual merits.